OpenStudy (anonymous):

Write 6.86 as a mixed number in simplest form.

jimthompson5910 (jim_thompson5910):

The second 6 (on the right) is in the hundredths place so that means that 0.86 = 86/100 and that means 6.86 = 6 + 0.86 = 6 + 86/100 make sure you reduce 86/100 and tell me what you get